OpenStudy (anonymous):

Look at this linear equation. 3x + 5y = 30 what equations can be used to solve for y in the linear equation above?

OpenStudy (anonymous):

So are we just supposed to look at it.

OpenStudy (anonymous):

idk im dumb lol

OpenStudy (anonymous):

help with mine plz

OpenStudy (anonymous):

i know how to solve this

OpenStudy (anonymous):

you have to work your way through the problem to get y by itself before you can solve the problem fully

OpenStudy (anonymous):

3x+5y=30 5y=3x+30 y= 3/5x+6

OpenStudy (anonymous):

whats the rest of the proble

OpenStudy (anonymous):

*problem

OpenStudy (anonymous):

thats all there was :/

OpenStudy (anonymous):

don't forget to change the sign when moving across the equal sign Ashley

OpenStudy (anonymous):

help me plz

OpenStudy (anonymous):

3x + 5y = 30 --- subtract 3x from both sides 5y = -3x + 30 -- divide both sides by 5 y = -3/5x + 6

OpenStudy (anonymous):

I suppose the equation to use to solvr for y is the slope intercept equation, which is y = mx + b

OpenStudy (anonymous):

typo....solve

OpenStudy (anonymous):

can u plz check my problem

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):

@uzair1928

OpenStudy (anonymous):

@uzair1928 bump your question

OpenStudy (anonymous):

I have to wait 15 min

OpenStudy (anonymous):

can i post it here

OpenStudy (anonymous):

you can post a link to it

OpenStudy (anonymous):

go for it

OpenStudy (anonymous):

Help please Graph a triangle (ABC) and reflect it over the x-axis to create triangle A'B'C'. Part 1: Describe the transformation using words. Make sure you refer to the characteristics and the coordinates. Part 2: Draw a line segment from point A to the reflecting line, and then draw a line segment from point A' to the reflecting line. What do you notice about the two line segments you drew? Part 3: Do you think you would see the same characteristic if you drew the line segment connecting B with the reflecting line and then B' with the reflecting line? How do you know?